OpenStudy (anonymous):

Let y = (ln((x^2)-1)^-3)/ 2. Find derivative

OpenStudy (anonymous):

y = (ln((x^2)-1)^-3)/ 2 dy/dx=d{(ln((x^2)-1)^-3)/ 2}/dx =d{-3ln((x^2)-1)/2}/d((x^2)-1) * d((x^2)-1)/dx

OpenStudy (anonymous):

Now, can u?